OpenStudy (anonymous):

find the zeros of this function. state the multiplicity of multiple zeros y=(x=2)^2(x-5)^4

OpenStudy (anonymous):

I'm assuming that the second = is actually a + 0 = (x+2)^2; x = -2, multiplicity of 2 0 = (x-5)^4; x = 5, multiplicity of 4 The multiplicity is the recurrence of that zero, in this case, the exponent. The zero is the value of x for which y would equal 0.